x, and S1 are the sample mean and sample variance from a population with mean μ| and variance ơf. Similarly, X2 and S1 are the sample mean and sample variance from a second population with mean μ and variance σ2. Assume that these two populations are independent, and the sample sizes from each population are n,and n2, respectively. (a) Show that X1-X2 is an unbiased estimator of μ1-μ2. (b) Find the standard error of X, -X. How could you estimate the standard error? (c) Show that the sample variance S1 defined by n 11 Ση11(Xi-X1)2 is an unbiased estimator of true 2 variance σ (d) Suppose that both populations have the same variance, that is, σ-σ: σ2 . Show that mi + n2-2 is an unbiased estimator of σ2
Answer #1

solution : oiven that unbiased estimat0918 Mer H2 are zi, 7, respecthere we know that unbiased estimators of 0,2,0,2 are S;?,(br. The standaard error of Xi - X2 mane Xn0 CH₂ , 021ni X2 und CM2 , 02² ingo X, - Xq wo CH - H2, of +02² Ines standand deviexpe=c727H 13 → EX, EH, v Că, 2 =VC Exi) = 11n? EUCxí) =o2 in, E (22) = VCX, ) + E CXj2 Det = colin,)th, ELX, 2) = (0,2 niothES? = I, C Cor+M? 2-n, (of in the = Ź, (2,0,2-0,2 f970 m equations and ESP = (no) infin ne in 2 (d). Both populations have th- cn,-1) E {S2 + (ne- E [S22] nit12-2 since 0,2-0,2-02 = (1,-1262+(12-1)002 - = 02 nitna-2 It is an onblased estimatoon , -X-